OpenStudy (anonymous):

pressure P is directly proportional to temperature T, P is 20 lb per square inch when T is 60 degrees Kelvin. What is the pressure when the temperature is 74 degrees kelvin?

OpenStudy (anonymous):

\[\frac{ 20 }{ 60 } = \frac{ x }{ 72 }\]

OpenStudy (anonymous):

err 74 im sorry.

OpenStudy (anonymous):

someone said 60/20=x/74

OpenStudy (anonymous):

Yeah thats the same thing. to solve you can cross multiply. and get 60x = 1440

OpenStudy (anonymous):

lets try your answer 24

OpenStudy (anonymous):

Thank You:) now hold up a little:)

OpenStudy (anonymous):

sorry it didnt work....im using wiley plus and it marked it wrong:(

OpenStudy (anonymous):

That cant be right...

OpenStudy (anonymous):

24.667 haha sorry

OpenStudy (anonymous):

let me try that...sorry i bug too much heheh

OpenStudy (anonymous):

miracle...THANK YOU! It worked lol

OpenStudy (anonymous):

lol my bad i used 72

OpenStudy (anonymous):

its ok thank you..are you a math tutor?

OpenStudy (anonymous):

N

OpenStudy (anonymous):

No haha, just an engineering student.

OpenStudy (anonymous):

wow, that's really good, difficult career but your smart...im undeclared, math is not my strength

OpenStudy (anonymous):

haha no worries, its a struggle all the time. just slow and steady

OpenStudy (anonymous):

That's true...I got another one fore you if that's ok with you. I need to construct an equation so that I can solve the problem. Earnings E are directly proportional to the time T worked, and E is $45 when T is 3hours. How long has a person worked if she earned $455?